OpenStudy (anonymous):

6. A hyperbola centered at (0, 0) has vertices (0, ±6) and one focus (0, −10). What is the standard form of the equation of the hyperbola? (1 point)

OpenStudy (anonymous):

@jhonyy9 @terenzreignz

OpenStudy (mertsj):

@ freshkid944 You're not here.

OpenStudy (anonymous):

im here

OpenStudy (mertsj):

The vertices are (0,6) and (0,-6) That tells us that the hyperbola is like this: 

OpenStudy (mertsj):

Therefore we know the following: The center is (0,0) and the value of "a" is 6

OpenStudy (mertsj):

Furthermore we are told that the foci are (0,10) and (0,-10) so we know that the value of "c" is 10

OpenStudy (mertsj):

If we have been paying attention in class, we know that for a hyperbola, a^2+b^2=c^2 and so we can calculate the value of b^2

OpenStudy (mertsj):

Once we know b^2 and a^2 we can plug those values into the standard equation of a hyperbola such as this one which is: \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]

OpenStudy (anonymous):

y squared over one hundred minus x squared over thirty-six equals one x squared over one hundred minus y squared over thirty-six equals one x squared over thirty-six minus y squared over sixty-four equals one y squared over thirty-six minus x squared over sixty-four equals one these are the answer choices @mertsj

OpenStudy (anonymous):

@jhonyy9

OpenStudy (mertsj):

What did you get for a and b?

OpenStudy (anonymous):

x squared over one hundred minus y squared over thirty-six equals one